# Towards Distillation-Resistant Large Language Models: An Information-Theoretic Perspective

### v1

@@ Abstract
Proprietary large language models (LLMs) embody substantial economic value and are generally exposed only as black-box APIs, yet adversaries can still exploit their outputs to extract knowledge via distillation. Existing defenses focus exclusively on text-based distillation, leaving the important logit-based distillation largely unexplored. In this work, we analyze this problem and present an effective solution from an information-theoretic perspective. We characterize distillation-relevant information in teacher outputs using the conditional mutual information (CMI) between teacher logits and input queries conditioned on ground-truth labels. This quantity captures contextual information beneficial for model extraction, motivating us to defend distillation via CMI minimization. Guided by our theoretical analysis, we propose learning a transformation matrix that purifies the original outputs to enhance distillation resistance. We further derive a CMI-inspired anti-distillation objective to optimize this transformation, which effectively removes distillation-relevant information while preserving output utility. Extensive experiments across multiple LLMs and strong distillation algorithms demonstrate that the proposed method significantly degrades distillation performance while preserving task accuracy, effectively protecting models’ intellectual property.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated remarkable capabilities and have been deployed in a variety of practical applications, such as multi-agent systems (Wu et al., 2024; Du et al., 2024). Training such models demands substantial resources, including large-scale datasets, extensive computational budgets, and significant human effort, rendering them highly valuable intellectual property.
@@
To safeguard their interests, commercial service providers typically isolate their proprietary models and only expose them through APIs. However, malicious attackers can exploit the model output texts or logits to perform knowledge distillation (KD) (Hinton et al., 2015; Kim and Rush, 2016; Gu et al., 2024), thus cheaply extracting model capabilities without any access to the weights or architecture.
To defend against such theft, early work has proposed LLM watermarking (Kirchenbauer et al., 2023; Ye et al., 2025; Liang et al., 2026) as a mitigation strategy, yet they offer only post-hoc verification and cannot directly prevent the extraction of proprietary knowledge. More recent efforts attempt to perturb the token distribution and modify model outputs to hinder KD (Savani et al., 2025; Li et al., 2025). However, these methods focus exclusively on text-based distillation, leaving the more widely discussed logit-based distillation largely unaddressed. Notably, a model’s output logits reveal the full probability distribution over tokens, conveying substantially richer information than the sampled token labels. Consequently, defending against such distillation is particularly challenging. This difficulty is further exacerbated by a growing body of work (Wang et al., 2021; Gu et al., 2024; Wang et al., 2025) that strengthens logit-based KD by more fully exploiting the rich information in these distributions. To address this challenge, we introduce an information-theoretic measure to quantify the distillation-relevant information in teacher logits and propose a transformation strategy to mitigate knowledge leakage.
@@
Fundamentally, LLMs perform a classification task to predict the next token for an ongoing sequence. The information in an output vector can be interpreted in two aspects: (1) the predicted label for the current input (i.e., the token corresponding to the maximum logit), and (2) additional contextual information beyond the label (Ye et al., 2024). Although different prefixes can yield the same next predicted token, their logits differ substantially due to varying contexts in semantics and syntax. Particularly, logit-based KD exploits such contextual information to transfer richer knowledge to the student model (Gu et al., 2024).
@@
To quantify the supervisory guidance from teacher outputs during distillation, we introduce conditional mutual information (CMI) as a principled measure of the conveyed information. Formally, let $X,Y,Z$ denote the input sequence, the ground-truth label, and the teacher logits for the given $X$, respectively. Then $\mathcal{I}(X;Z|Y)$ measures how much information $Z$ provides about $X$ conditioned on $Y$. This precisely aligns with the paradigm of logit-based KD that exploits information encoded in teacher logits beyond the label to enhance student performance. Therefore, a natural defense is to purify the teacher logits to minimize the conditional mutual dependence between $X$ and $Z$. To achieve this goal, we propose learning a transition matrix $\mathcal{M}$ that post-processes the output logits (see Figure 1), which we theoretically validate to reduce $\mathcal{I}(X;Z|Y)$. To effectively learn this matrix and decrease CMI, we further expand $\mathcal{I}(X;Z|Y)$ and decompose the minimization into two sub-problems based on the Information Bottleneck (IB) principle (IB开盒): (1) increasing the dependency between the logits $Z$ and the ground-truth label $Y$, and (2) weakening the correlation between the input $X$ and the logits $Z$. Guided by this analysis, we devise an optimization objective with two corresponding loss terms. We first derive a theoretical upper bound for the first subproblem, which formally justifies that minimizing the cross-entropy between $Z$ and $Y$ is a practical surrogate objective. The latter sub-task aims to reduce the relevance of $X$ and $Z$, yet directly computing or estimating this term for LLMs is intractable. To circumvent this difficulty, we note that under the KD scenario, the information in $Z$ about $X$ is ultimately leveraged through gradients propagated to the student during optimization. In this sense, these gradients serve as the practical carrier of the guidance provided by teacher logits. Building on this insight, we employ a surrogate student model to simulate distillation and maximize the angular deviation between the gradients induced by the transformed logits $Z^{\prime}$ and those from the original $Z$. By iteratively optimizing the transformation $\mathcal{M}$ with the proposed loss function, we effectively reduce the distillation-beneficial contextual information in teacher outputs, thereby strengthening their resistance against knowledge extraction attacks.
@@
  • To the best of our knowledge, we are the first to defend LLMs against logit-based KD. We analyze the problem from an information-theoretic perspective and propose learning a matrix that post-processes output logits to minimize their CMI with the input.
  • Guided by theoretical analysis, we design two complementary loss functions to optimize the transformation matrix, which effectively removes the contextual information within the logits that could be exploited by attackers while preserving output utility.
  • Extensive experiments across various LLMs on three widely used datasets against four strong distillation algorithms demonstrate the outstanding defense performance of the proposed method.
@@ 2.1 Knowledge Distillation for Large Language Models
Knowledge Distillation (KD) is a widely adopted paradigm for transferring knowledge from high-capacity teacher models to smaller and less capable student models. For large language models (LLMs), KD is usually realized through two paradigms: sequence-level (text-based) distillation and logit-based (response-level) distillation.
Sequence-level knowledge distillation trains the student model using texts generated by the teacher as hard-label supervision, treating the teacher-produced outputs as pseudo-labels and optimizing the student via standard maximum-likelihood estimation (Kim and Rush, 2016). In contrast, besides the SFT loss based on ground-truth labels, logit-based KD additionally uses the teacher’s soft output distributions to distill fine-grained predictive information by aligning the teacher and student distributions. The pioneering logit-based KD method (Hinton et al., 2015) uses the Kullback-Leibler divergence (KLD) to measure the discrepancy between the student and the teacher distributions. Reverse KLD (RKLD) employs a reverse version of KL divergence by changing the reference distribution in KL from the teacher to the student, allowing the student to concentrate its limited capacity on the teacher’s high-probability regions (Gu et al., 2024). AlphaNet (Wang et al., 2021) introduces $\alpha$-divergence, providing a tunable measure between KLD and RKLD to penalize over-estimation and under-estimation in KD. ABKD (Wang et al., 2025) further proposes a novel $\alpha\text{-}\beta$-divergence to balance hardness-concentration and confidence-concentration, generalizing FKLD and RKLD to a broader family of divergences.
@@ 2.2 Protecting LLMs from Knowledge Distillation
The great success of KD poses a significant threat to the protection of intellectual property in LLMs. To mitigate this issue, watermarking (Kirchenbauer et al., 2023; Liang et al., 2026) seeks to embed an identifiable and verifiable signature into a model’s behavior or parameters, enabling the model owner to verify their ownership. Meanwhile, Fingerprinting (Ye et al., 2025; Xu et al., 2024) assigns a unique and traceable identity to each distributed model instance, enabling the model owner to attribute leaked or misused models to their source. Despite their usefulness, both approaches are inherently post hoc, indicating that they facilitate the identification of model stealing after it has occurred, but cannot prevent the extraction process itself.
Another research line attempts to directly solve this threat by modifying the token distribution during decoding to obtain generated texts that are resilient to distillation. The pioneering work (Savani et al., 2025) proposes an anti-distillation decoding strategy that actively poisons reasoning traces by injecting a learned perturbation into the token-level probability distribution. Concretely, a proxy student first computes gradients on a small hold-out set of teacher-generated sequences, yielding a fixed poisoning direction $\nabla\ell$. During inference, the teacher performs two forward passes for each candidate token with proxy-student parameters perturbed by $\pm\eta\nabla\ell$, and estimates the resulting finite-difference logit shift. This shift is then integrated into the teacher’s logits to form an adjusted sampling distribution. Sampling from this modified distribution implicitly steers any downstream student trained on these traces toward higher loss, thereby degrading the effectiveness of distillation. DOGe (Li et al., 2025) adopts a more direct approach that maximizes the divergence between teacher and student output distributions to make the teacher’s predictions harder to imitate. Specifically, DOGe fine-tunes the teacher’s final linear layer via the negative KL-divergence between the temperature-scaled softmax of the teacher logits and the detached softmax of a proxy student, enlarging the teacher–student discrepancy.
@@ 3.1 Notation
Let $\mathcal{V}$ be the vocabulary and $|\mathcal{V}|<\infty$. Let $X\in\mathcal{V}^{*}$ be the input sequence consisting of the user’s query and the ongoing generated response, and $Y\in\mathcal{V}$ be the ground-truth label, i.e., the next token for the current input $X$.
Let $f:\mathcal{V}^{*}\to\mathbb{R}^{|\mathcal{V}|}$ be the large language model, and $Z=f(X)$ be the logit vector for current input $X$. The next token prediction is sampled from the distribution $\sigma(Z)$, where $\sigma(\cdot)$ denotes the softmax function.
@@ 3.2 Knowledge Distillation
where $l$ denotes the length of the ground-truth answer $o$, and $\text{Dist}(\cdot\|\cdot)$ denotes the distance metric such as Kullback–Leibler divergence (KLD). In general, the objective of logit-based distillation algorithms can be formulated as:
@@
where $\alpha$ is a factor set to $0.5$ by default, and $\mathcal{L}_{\mathrm{SFT}}$ denotes the supervised fine-tuning (SFT) loss, i.e., the negative log-likelihood of the predictive probability of the ground-truth label. Since the logits convey more contextual information than the predictive label alone, KD typically outperforms standard SFT on many downstream tasks. Based on this formulation, various improved algorithms have been proposed, including novel designs of distance metrics such as reverse KLD (Gu et al., 2024) and AB-KLD (Wang et al., 2025).
@@ 3.3 Information Bottleneck
The Information Bottleneck (IB) theory (IB提出; IB深度; IB开盒) highlights that a well-trained model should compress redundant information from inputs while preserving key information for tasks. The IB objective is formulated as:
@@
Figure 2: An illustration of the proposed defense framework. As shown in the left subfigure, we introduce a surrogate student model as an adversary and simulate the KD process. We compute the KD loss $\mathcal{L}_{\text{total}}$ and extract its gradients $\bm{g}$ with respect to the student parameters to quantify the guidance signal carried by the teacher outputs. The right subfigure explains the loss computation used for training the transformation matrix. To diminish context information in teacher logits, we design to minimize the KLD between the transformed logits $Z^{\prime}$ and the ground-truth label $Y$, while simultaneously maximizing the angular deviation between the gradients $\bm{g}$ and $\bm{g}^{\prime}$ calculated by the original and transformed logits, respectively.
@@ 4.1 Conditional Mutual Information-Based Defense
The CMI $\mathcal{I}(X;Z|Y)$ measures how much information $Z$ reveals about $X$ given the ground-truth label $Y$. The additional contextual information that $Z$ carries beyond the label aligns exactly with the goal of logit-based distillation, as shown in Equation 1. Therefore, we propose to defend teacher knowledge by minimizing $\mathcal{I}(X;Z|Y)$, preventing such contextual information from being extracted. This objective can also be interpreted through entropy and uncertainty. Specifically, $\mathcal{I}(X;Z|Y)=\mathcal{H}(X|Y)-\mathcal{H}(X|Z,Y)$, which quantifies the reduction in uncertainty of $X$ when $Z$ is additionally observed. For a fixed training dataset, $\mathcal{H}(X|Y)$ remains constant. Hence, minimizing $\mathcal{I}(X;Z|Y)$ is equivalent to maximizing $\mathcal{H}(X|Z,Y)$, i.e., making $Z$ contain as little information about $X$ as possible to increase the uncertainty about $X$ given the label $Y$.
@@
As an intuitive extreme case, if $Z$ ideally degenerates to the one-hot label $Y$, i.e., it contains no contextual information beyond the category, then logit-based knowledge distillation collapses to standard label-only supervised fine-tuning. In this case, the CMI attains its minimum value $\mathcal{I}(X;Z|Y)=\mathcal{I}(X;Y|Y)=0$.
To reduce the contextual information in teacher logits, we propose a lightweight post-processing strategy that further modifies output logits while keeping the teacher completely frozen. Concretely, we introduce a learnable matrix $\mathcal{M}$ to transform the original teacher output into a refined one. Before delving into the full training pipeline, we show that such a transformation admits an intriguing information-theoretic property that formally guarantees that the refined logits become more context-compressed:
@@
$\mathcal{I}(X;Z|Y)\geq\mathcal{I}(X;Z^{\prime}|Y).$

### v2

@@ Abstract
Proprietary large language models (LLMs) embody substantial economic value and are generally exposed only as black-box APIs, yet adversaries can still exploit their outputs to extract knowledge via distillation. Existing defenses focus exclusively on text-based distillation, leaving the important logit-based distillation largely unexplored. In this work, we analyze this problem from an information-theoretic perspective and present an effective solution. Specifically, we characterize distillation-relevant information in teacher outputs via the conditional mutual information (CMI) between teacher logits and input queries, conditioned on ground-truth labels. This quantity captures contextual information that facilitates knowledge extraction, motivating us to defend distillation via CMI minimization. Guided by our theoretical analysis, we propose learning a transformation matrix that purifies the original outputs to enhance distillation resistance. We further derive a CMI-inspired anti-distillation objective to optimize this transformation, which effectively reduces distillation-relevant information while preserving output utility. Extensive experiments across multiple LLMs and strong distillation algorithms demonstrate that the proposed method significantly degrades distillation performance while preserving task accuracy, effectively protecting models’ intellectual property.
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated remarkable capabilities and have been deployed in a variety of practical scenarios, such as healthcare applications Goyal et al. (2024) and multi-agent systems Wu et al. (2024); Du et al. (2024). Training such huge models demands substantial resources, including large-scale datasets, extensive computational budgets, and significant human effort, rendering them highly valuable intellectual property. To safeguard their interests, commercial service providers typically isolate their proprietary models and only expose them through APIs. However, malicious attackers can exploit the output texts or logits to perform knowledge distillation (KD) Hinton et al. (2015); Kim and Rush (2016); Gu et al. (2024), thus cheaply extracting model capabilities without any access to the weights or architecture.
To defend against such theft, previous work has explored a range of LLM watermarking Kirchenbauer et al. (2023); Ye et al. (2025); Liang et al. (2026) and fingerprinting techniques Chen et al. (2023); Yu et al. (2021); Peng et al. (2022) for ownership verification, yet they provide only post-hoc verification and cannot directly prevent the extraction of proprietary knowledge. More recent efforts attempt to perturb the token distribution and modify model outputs to hinder KD Savani et al. (2025); Li et al. (2025). Despite the success, these methods focus exclusively on text-based distillation, leaving the more widely discussed logit-based distillation largely unaddressed. Notably, a model’s output logits reveal the full probability distribution over tokens, conveying substantially richer information than the sampled token labels. Consequently, defending against such distillation is particularly more challenging. This difficulty is further exacerbated by a growing body of work Wang et al. (2021); Gu et al. (2024); Wang et al. (2025) that strengthens logit-based KD by more fully exploiting the rich information in these distributions.
This work investigates this critical challenge by introducing an information-theoretic measure to quantify the distillation-relevant information within teacher logits, based on which we further propose an output transformation strategy to mitigate knowledge leakage.
@@
Fundamentally, LLMs perform a classification task to predict the next token for an ongoing input sequence. The information in an output vector can be interpreted in two aspects: (1) the predicted label for the current input (i.e., the token corresponding to the maximum logit), and (2) additional contextual information beyond the label Ye et al. (2024). Although different prefixes can yield the same next predicted token, their logits differ substantially due to varying contexts in semantics and syntax. Particularly, logit-based KD exploits such contextual information to transfer richer knowledge to the student model Gu et al. (2024).
@@
To quantify the supervisory guidance from teacher outputs during distillation, we introduce conditional mutual information (CMI) as a principled measure of the conveyed information. Formally, let $X,Y,Z$ denote the input sequence, the ground-truth label, and the teacher logits for the given $X$, respectively. Then $\mathcal{I}(X;Z|Y)$ measures how much information $Z$ provides about $X$ conditioned on $Y$. This precisely aligns with the paradigm of logit-based KD that exploits information encoded in teacher logits beyond the label to enhance student performance. Therefore, a natural defense is to purify the teacher logits to minimize the conditional mutual dependence between $X$ and $Z$. To achieve this goal, we propose learning a transformation matrix $\mathcal{M}$ that post-processes the output logits (see Figure 1), which we theoretically validate to reduce $\mathcal{I}(X;Z|Y)$. To effectively learn this matrix and decrease CMI, we further expand $\mathcal{I}(X;Z|Y)$ and decompose the minimization into two sub-problems based on the Information Bottleneck (IB) principle IB开盒: (1) increasing the dependency between the logits $Z$ and the ground-truth label $Y$, and (2) weakening the correlation between the input $X$ and the logits $Z$. Guided by this analysis, we devise an optimization objective with two corresponding loss terms. We first derive a theoretical upper bound for the first subproblem, which formally justifies that minimizing the cross-entropy between $Z$ and $Y$ is a practical surrogate objective. The latter sub-task aims to reduce the relevance of $X$ and $Z$, yet directly computing or estimating this term for LLMs is intractable. To circumvent this difficulty, we note that under the KD scenario, the information in $Z$ about $X$ is ultimately leveraged through gradients propagated to the student during optimization. In this sense, these gradients serve as the practical carrier of the guidance provided by teacher logits. Building on this insight, we employ a surrogate student model to simulate distillation and maximize the angular deviation between the gradients induced by the transformed logits $Z^{\prime}$ and those from the original $Z$. By iteratively optimizing the transformation $\mathcal{M}$ with the proposed loss function, we effectively reduce the distillation-beneficial contextual information in teacher outputs, thereby strengthening their resistance against knowledge extraction attacks.
@@
  • To the best of our knowledge, we are the first to defend LLMs against logit-based KD. We analyze the problem from an information-theoretic perspective and propose learning a matrix that post-processes output logits to minimize their CMI with the input query.
  • Guided by theoretical analysis, we design two complementary loss functions to optimize the transformation matrix, which effectively removes the contextual information within teacher logits that could be exploited by malicious attackers while preserving output utility.
  • Extensive experiments across multiple LLMs such as Llama 3 series across three widely used datasets against four strong distillation algorithms demonstrate the outstanding defense performance of the proposed method in safeguarding model intellectual property.
@@ 2.1 Knowledge Distillation for Large Language Models
Knowledge Distillation (KD) is a widely adopted paradigm for transferring useful knowledge from more powerful teacher models to less capable (generally smaller) student models. For large language models (LLMs), the distillation algorithm is typically achieved through two paradigms: sequence-level (text-based) and logit-based (response-level) approaches.
Sequence-level KD trains the student model using teacher-generated texts as hard-label supervision, treating teacher-produced outputs as pseudo-labels and optimizing the student via standard maximum-likelihood estimation Kim and Rush (2016). In addition to the SFT loss based on ground-truth labels, logit-based KD utilizes the teacher’s soft output distributions to distill fine-grained predictive information by aligning the teacher and student distributions. The pioneering logit-based KD method Hinton et al. (2015) uses Kullback-Leibler divergence (KLD) to measure the discrepancy between student and teacher distributions. Reverse KLD (RKLD) employs a reverse version of KL divergence by changing the reference distribution in KL from the teacher to the student, allowing the student to concentrate its limited capacity on the teacher’s high-probability regions Gu et al. (2024). AlphaNet Wang et al. (2021) introduces $\alpha$-divergence, providing a tunable measure between KLD and RKLD to penalize over-estimation and under-estimation in KD. ABKD Wang et al. (2025) further proposes a novel $\alpha\text{-}\beta$-divergence to balance hardness-concentration and confidence-concentration, generalizing FKLD and RKLD to a broader family of divergences.
@@ 2.2 Protecting LLMs from Knowledge Distillation
The success of KD poses a significant threat to the protection of intellectual property in LLMs. To mitigate this issue, watermarking Kirchenbauer et al. (2023); Liang et al. (2026) embeds an identifiable and verifiable signature into a model’s behavior or parameters to verify the ownership. Fingerprinting Ye et al. (2025); Xu et al. (2024) assigns a unique and traceable identity to each distributed model instance, enabling the model owner to attribute leaked or misused models. However, both approaches are inherently post hoc, i.e., they enable the identification of model stealing after it has occurred, but cannot prevent the extraction process itself.
Another research line aims to directly remove this threat by modifying the decoding distribution to obtain generated texts that are resilient to distillation. The pioneering work Savani et al. (2025) proposes an anti-distillation decoding strategy that actively poisons reasoning traces by injecting a learned perturbation into the token-level probability distribution. Concretely, a proxy student first computes gradients on a small hold-out set of teacher-generated sequences, yielding a fixed poisoning direction $\nabla\ell$. During inference, the teacher performs two forward passes for each candidate token with proxy-student parameters perturbed by $\pm\eta\nabla\ell$, and estimates the resulting finite-difference logit shift. This shift is then integrated into the teacher’s logits to form an adjusted sampling distribution. Sampling from this modified distribution implicitly steers any downstream student trained on these traces toward higher loss, thereby degrading the effectiveness of distillation. DOGe Li et al. (2025) adopts a more direct approach that maximizes the divergence between teacher and student output distributions to make the teacher’s predictions harder to imitate. Specifically, DOGe fine-tunes the teacher’s final linear layer with an adversarial loss defined as the negative KL-divergence between the temperature-scaled softmax of the teacher logits and the detached softmax of a proxy student, enlarging the teacher–student discrepancy.
@@ 3.1 Notation
Let $\mathcal{V}$ be the vocabulary and $|\mathcal{V}|<\infty$. Let $X\in\mathcal{V}^{*}$ be the input sequence consisting of the user’s query and the ongoing generated response, and let $Y\in\mathcal{V}$ be the ground-truth label, i.e., the next token for the current input $X$. We define the large language model as a function $f:\mathcal{V}^{*}\to\mathbb{R}^{|\mathcal{V}|}$, and let $Z=f(X)$ be the resulting logit vector for input $X$. The next token prediction is sampled from the distribution $\sigma(Z)$, where $\sigma(\cdot)$ denotes the softmax function.
@@ 3.2 Knowledge Distillation
where $l$ denotes the length of the ground-truth answer $o$, and $\text{Dist}(\cdot\|\cdot)$ denotes the distance metric such as Kullback–Leibler divergence (KLD). In general, the optimization objective of logit-based distillation algorithms can be formulated as:
@@
where the factor $\alpha=0.5$ by default, and $\mathcal{L}_{\mathrm{SFT}}$ denotes the supervised fine-tuning (SFT) loss, i.e., the negative log-likelihood of the predictive probability of the ground-truth label. Since the logits carry more contextual information than the predictive label alone, KD typically outperforms standard SFT on many downstream tasks. Based on this formulation, various improved algorithms have been proposed, including novel designs of distance metrics such as reverse KLD Gu et al. (2024) and AB-KLD Wang et al. (2025).
@@ 3.3 Information Bottleneck
The Information Bottleneck (IB) theory IB提出; IB深度; IB开盒 highlights that a well-trained model should compress redundant information from inputs while preserving key information for tasks. The IB objective is formulated as:
@@ 4.1 Conditional Mutual Information-Based Defense
The CMI $\mathcal{I}(X;Z|Y)$ measures how much information $Z$ reveals about $X$ given the ground-truth label $Y$. The additional contextual information that $Z$ carries beyond the label aligns exactly with the goal of logit-based distillation, as revealed by Eq. (1) and (2). Therefore, we propose to protect teacher knowledge by minimizing $\mathcal{I}(X;Z|Y)$, i.e., suppressing the contextual information encoded in teacher logits. This objective can also be interpreted through entropy and uncertainty. Specifically, $\mathcal{I}(X;Z|Y)=\mathcal{H}(X|Y)-\mathcal{H}(X|Z,Y)$, which quantifies the reduction in uncertainty of $X$ when $Z$ is additionally observed. For a fixed training dataset, $\mathcal{H}(X|Y)$ remains constant. Hence, minimizing $\mathcal{I}(X;Z|Y)$ is equivalent to maximizing $\mathcal{H}(X|Z,Y)$, i.e., making $Z$ contain as little information about $X$ as possible to increase the uncertainty about $X$ given the label $Y$.
@@
As an intuitive extreme case, if $Z$ ideally degenerates to the one-hot label $Y$, i.e., it contains no contextual information beyond the category, then logit-based knowledge distillation collapses to standard label-only supervised fine-tuning. In this case, the CMI indeed attains its minimum value $\mathcal{I}(X;Z|Y)=\mathcal{I}(X;Y|Y)=0$.
To achieve the goal of CMI minimization, we propose a lightweight post-processing strategy that further modifies output logits while keeping the teacher model completely frozen. Concretely, we introduce a learnable matrix transformation $\mathcal{M}(\cdot)$ that maps the original teacher output to refined ones. Before delving into the full training pipeline, we show that such a transformation admits an intriguing information-theoretic property, which formally guarantees that the refined logits become more effectively compressed in terms of contextual information:
@@
$\mathcal{I}(X;Z|Y)\geq\mathcal{I}(X;Z^{\prime}|Y).$ (5)
Figure 2: An illustration of the proposed method. In the left subfigure, we introduce a surrogate student model as an adversary and simulate the KD process. We compute the gradient $\bm{g}$ of the distillation loss $\mathcal{L}_{\text{total}}$ with respect to the student parameters to quantify the guidance signal carried by teacher outputs. The right subfigure explains the training loss for the transformation matrix. To diminish context information in teacher logits, we minimize the cross-entropy between the transformed logits $Z^{\prime}$ and the ground-truth label $Y$, while simultaneously maximizing the angular deviation between the gradients $\bm{g}$ and $\bm{g}^{\prime}$ calculated by the original and transformed logits, respectively.